OpenStudy (anonymous):

write a point slope form of the line through point J(4,1) with slope -4

OpenStudy (anonymous):

write an equation point slope form of the line through point J(4,1) with slope -4 thats what i ment

OpenStudy (amistre64):

y-Jy=m(x-Jx) is the form

OpenStudy (anonymous):

We can use the point-slope formula y - y1 = m(x - x1) Where m = -4 x1 = 4 y1 =1 and we have y - 1 = -4(x - 4) Can we leave the answer in this format?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

you know that the equation should be y = m*x + b , u have the point and the slope to find the equation of the line passing through that point sub every thing in and solve for b and u will have your answer